OpenStudy (anonymous):

Dr. Gupta is considering a 30 year mortgage at 9% interest. She can make payments of $2500 a month. What size loan can she afford? (Round to your nearest tenth) please help

OpenStudy (amistre64):

think of this as an annuity

OpenStudy (amistre64):

i loan follows the pattern: \[B_n=B_0(k)^n-P(\frac{1-k^n}{1-k})~:~k=1+\frac{r}{12}, n=12*years\]

OpenStudy (amistre64):

at B_n the loan is paid off = 0 at B_o the loan amount is taken out; sove for B_o